# Structure-Preserving Oscillation-Eliminating Discontinuous Galerkin Schemes for Ideal MHD Equations: Locally Divergence-Free and Positivity-Preserving

### v1

@@ 4.2.5 MHD blast wave
We now run the blast wave problem to test the ability of our scheme in cases of strong shocks and rarefaction wave. In the blast wave problem, the over-pressured region, located at the center of strongly magnetized medium with low plasma $\beta$, will generate an MHD blast wave, which drives the fast outward magnetic shocks to compress the plasma and magnetic field ahead, and generates rarefaction wave internally. For this test, we adopted the version of Stone et al. [66] with the initial settings as follows:
$\begin{array}[]{l}\rho=1,\quad u_{1}=u_{2}=u_{3}=0,\quad B_{1}=B_{2}=1/\sqrt{2},\quad B_{3}=0,\quad p=\left\{\begin{aligned} &10,&\text{ if }\sqrt{(x-1.5)^{2}+y^{2}}<0.1,\\ &0.1,&\text{ otherwise }.\\ \end{aligned}\right.\end{array}$
We use $\gamma=5/3$ and fix the boundary conditions. The simulation is run on the domain $\left[1,2\right]\times\left[-0.5,0.5\right]$ until the final time of $t=0.2$.
In Figure 12, we show the density, thermal pressure, velocity, and magnetic pressure for the third-order PP LDF OEDG scheme at time $t=0.2$ with $400\times 400$ cells. The results are consistent with those in [59, 60]. One can see that a circular blast wave propagates outward, and a rarefaction wave propagates inward. The symmetry of the problem is well preserved. Our scheme captures sharp shocks and remains smooth post-shock regions, which demonstrates the robustness of our method in highly magnetized shock configurations.
Figure 12: Maps of (a) density $\rho$, (b) thermal pressure $p$, (c) velocity $|\boldsymbol{u}|$, and (d) magnetic pressure $|\boldsymbol{B}|^{2}/2$ at time $t=0.2$ with $400\times 400$ cells for MHD blast wave problem.
@@ 4.2.6 MHD jet flows
The last example considers three challenging MHD jet problems, to further test the performance of our scheme in capturing shocks and maintaining positivity. MHD jet flows were proposed in Wu and Shu [43] by adding a magnetic field in the gas dynamical jet of Balsara [38]. Since there are strong shock waves, shear flows and interface instabilities in high Mach number jet with strong magnetic field and huge kinetic energy, negative pressure is very likely to be produced in the numerical simulations.
Our computational domain is $[-0.5,0.5]\times[0,1.5]$, and the adiabatic index $\gamma=1.4$. Initially, we set $(\rho,p)=(0.1\gamma,1)$ in the static ambient medium with a magnetic field $(0,B_{2},0)$. As observed in [43], the larger $B_{2}$ takes, the more challenging the test becomes. On the bottom boundary, a Mach 800 dense jet with $(\rho,p,u_{1},u_{2},u_{3})=(\gamma,1,0,800,0)$ is injected along the nozzle $\{y=0,|x|<0.05\}$. We use the fixed inflow condition on the nozzle, and outflow conditions on the other boundary conditions. For this problem, we set the computational domain as $[0,0.5]\times[0,1.5]$, i.e., take the right half of the whole domain, and utilize the reflecting boundary condition at $x=0$. The domain $[0,0.5]\times[0,1.5]$ is divided into $200\times 600$ cells. We first test our scheme with two magnetized cases: (i) $B_{2}=\sqrt{200}$; (ii) $B_{2}=\sqrt{2000}$. The numerical results of density logarithm $\log_{10}\rho$, thermal pressure logarithm $\log_{10}p$, and velocity $|\boldsymbol{u}|$ for two cases are displayed in Figure 13 and Figure 14, respectively. These numerical results clearly show the time evolution of flow structures, and the results agree with those in [43, 44]. Note that our scheme does a great job of capturing the Mach shock wave at the jet head and the beam/cocoon interface with high resolution. Specifically, the shear flows near the beam/cocoon interfaces of jets are clearly presented. In all the simulations, no negative pressure or negative density was observed. We also run our code on the test of higher Mach 10000 dense jet with stronger magnetized case of $B_{2}=\sqrt{20000}$. Figure 15 shows the time resolution of flow structures, including density logarithm $\log_{10}\rho$, thermal pressure logarithm $\log_{10}p$, and velocity $|\boldsymbol{u}|$. Our scheme exhibits good performance and robustness in such an extreme test. We found that if we turn off any of OE procedure, PP limiting procedure, and suitably discretized Godunov-Powell source term, the code will not be able to run jet tests with strong magnetic fields due to negative pressure produced in the cell averages of the numerical solution. This, again, demonstrates the importance of OE procedure, PP limiting procedure, and suitably discretized Godunov-Powell source term for PP property of the scheme.
@@
Figure 13: Maps of density logarithm (top), thermal pressure logarithm (middle), and velocity (bottom) for the Mach 800 jet problem with $B_{2}=\sqrt{200}$. From left to right: $t=0.001,0.0015,0.002$.
@@
Figure 14: Maps of density logarithm (top), thermal pressure logarithm (middle), and velocity (bottom) for the Mach 800 jet problem with $B_{2}=\sqrt{2000}$. From left to right: $t=0.001,0.0015,0.002$
Figure 15: Maps of density logarithm (top), thermal pressure logarithm (middle), and velocity (bottom) for the Mach 10000 jet problem with $B_{2}=\sqrt{20000}$. From left to right: $t=0.00005,0.0001,0.00015$
@@ 5 Concluding remarks
In this paper, we extend the OEDG method to ideal MHD equations, and propose an LDF OEDG scheme by using the LDF DG scheme as the basis scheme. In the LDF OEDG scheme, the OE procedures are performed on the Legendre basis and LDF basis respectively, as a result, the proposed scheme can effectively control the spurious oscillations while automatically maintaining the LDF property of the computed magnetic field. By alternately evolving each Runge-Kutta stage and an OE procedure, the implementation of the LDF OEDG scheme is effective and computational cheap on general meshes. Moreover, based on the PP limiter, HLL flux and proper discretized Godunov-Powell source term added to the evolution equations of the cell averages, we build a framework of PP LDF OEDG scheme on Cartesian meshes, and rigorously analyze the PP property of the proposed scheme. We derive the PP CFL condition on Cartesian meshes via the general convex decomposition of the cell averages, and also list the PP CFL conditions corresponding to the Zhang-Shu convex decomposition and optimal convex decomposition. Using third-order SSP Runge-Kutta time discretization, we test several MHD examples in one and two dimensions. The numerical results confirm that the proposed scheme achieves the expected optimal convergence order and behaves robust even in some tough MHD problems with low density or low pressure. Our next work is to conduct MHD numerical simulations based on the PP LDF OEDG schemes on unstructured meshes.

### v2

@@ 4.2.5 MHD blast wave
We now turn to a 2D blast wave test to further verify the capability of our structure-preserving OEDG scheme in handling strong shocks and rarefaction waves. This test involves an over-pressured region situated at the center of a strongly magnetized medium with low plasma $\beta$. The scenario leads to the formation of an MHD blast wave that drives fast outward magnetic shocks, compressing both the plasma and magnetic fields ahead, while internally generating a rarefaction wave. We utilize the setup in [69], with the initial conditions specified as
$\rho=1,~{}u_{1}=u_{2}=u_{3}=0,~{}B_{1}=B_{2}=\frac{1}{\sqrt{2}},~{}B_{3}=0,~{}p=\left\{\begin{aligned} &10,&\text{ if }\sqrt{(x-1.5)^{2}+y^{2}}<0.1,\\ &0.1,&\text{ otherwise}.\\ \end{aligned}\right.$
The adiabatic index $\gamma$ is set to $5/3$, and Dirichlet boundary conditions matching the initial values are applied. The simulation is conducted over the domain $\left[1,2\right]\times\left[-0.5,0.5\right]$ and runs until a final time of $t=0.2$.
Figure 13 presents the results at $t=0.2$ computed using our third-order PP LDF OEDG scheme on a $400\times 400$ mesh, showing the density, thermal pressure, velocity, and magnetic pressure. The results align well with those found in [70, 64]. The visualizations clearly illustrate that a circular blast wave propagates outward, while a rarefaction wave moves inward. The numerical solution maintains symmetry excellently. Our method successfully captures the shocks with high resolution and remains free from nonphysical oscillations, demonstrating its robustness and reliability in handling highly magnetized shock configurations. This performance underscores the capability of our structure-preserving OEDG scheme to effectively manage complex dynamic interactions in MHD environments.
Figure 13: The density $\rho$ (top left), thermal pressure $p$ (top right), velocity $\|\boldsymbol{u}\|$ (bottom left) and magnetic pressure $\|\boldsymbol{B}\|^{2}/2$ (bottom right) at time $t=0.2$ for the MHD blast wave problem with $400\times 400$ cells.
@@ 4.2.6 MHD jet flows
In our final example, we explore two challenging MHD jet problems to assess our scheme’s capability in capturing shocks and preserving positivity under demanding conditions. These MHD jet flows were proposed in [47] by adding a magnetic field in the gas dynamical jet of [41]. Since there are strong shock waves, shear flows and interface instabilities in high Mach number jet with strong magnetic field and huge kinetic energy, negative pressure is very likely to be produced in the numerical simulations.
Our computational domain is $[-0.5,0.5]\times[0,1.5]$, and the adiabatic index is $\gamma=1.4$. Initially, we set $(\rho,p)=(0.1\gamma,1)$ in the static ambient medium with a magnetic field $(0,B_{2},0)$. As observed in [47], the larger $B_{2}$ is set, the more challenging the test becomes. On the bottom boundary, a dense jet described by $(\rho,p,u_{1},u_{2},u_{3})=(\gamma,1,0,u_{2},0)$ is injected through the nozzle $\{y=0,|x|<0.05\}$, with the Mach number determined by $u_{2}$. We employ fixed inflow conditions at the nozzle and outflow conditions on other boundaries. Reflecting boundary conditions are used at $x=0$, with the domain limited to $[0,0.5]\times[0,1.5]$ on the right half, divided into $200\times 600$ cells.
First, we simulate a Mach 800 dense jet in a magnetized environment with $B_{2}=\sqrt{2000}$. The numerical results, showcasing density logarithm $\log_{10}\rho$, thermal pressure logarithm $\log_{10}p$, and velocity magnitude $\|\boldsymbol{u}\|$, are displayed in Figure 14. These results vividly depict the evolution of jet flow structures, consistent with those reported in [47, 48]. Our scheme accurately captures the Mach shock wave at the jet head and the beam/cocoon interface with high resolution. Specifically, the shear flows near the beam/cocoon interfaces of jets are clearly presented.
We also conduct a more demanding test with a Mach 10000 dense jet and a stronger magnetic field characterized by $B_{2}=\sqrt{20000}$. Figure 15 illustrates the time resolution of flow structures, including $\log_{10}\rho$, $\log_{10}p$, and $\|\boldsymbol{u}\|$. Throughout all simulations, no instances of negative pressure or density were observed, highlighting our scheme’s robustness even under extreme conditions. However, if we disable the LDF OE procedure, the PP limiting procedure, or the upwind discrete Godunov–Powell source term, the code immediately fails in these challenging tests due to negative pressure in the numerical solution. This underscores the essential roles of the LDF OE procedure, PP limiting procedure, and upwind discrete Godunov–Powell source terms in ensuring the PP property and overall robustness of our scheme.
@@
Figure 14: The density logarithm (top), thermal pressure logarithm (middle) and velocity (bottom) for the Mach 800 jet problem with $B_{2}=\sqrt{2000}$. From left to right: $t=0.001,0.0015,0.002$.
@@
Figure 15: The density logarithm (top), thermal pressure logarithm (middle), and velocity (bottom) for the Mach 10000 jet problem with $B_{2}=\sqrt{20000}$. From left to right: $t=0.00005,0.0001,0.00015$.
@@ 5 Concluding remarks
In this paper, we have proposed a structure-preserving, oscillation-eliminating discontinuous Galerkin (OEDG) method for ideal magnetohydrodynamics (MHD). This method stands out for its capability to effectively suppresses potential spurious oscillations in the DG solution while preserving the key physical structures, including the divergence-free constraint of magnetic field and the positivity of density and pressure. Based on a novel damping equation, the locally divergence-free (LDF) OE procedure is designed to eliminate spurious oscillations while automatically maintaining an LDF magnetic field. This procedure is implemented after each Runge–Kutta stage in a non-intrusive way, and thus can be easily integrated into existing DG codes as an independent module. Its integration into the LDF DG framework maintains beneficial properties such as conservation, local compactness, and optimal convergence rates. A rigorous positivity-preserving (PP) analysis has been conducted for the LDF OEDG method with the HLL flux. This analysis utilizes the geometric quasi-linearization (GQL) approach to transform the nonlinear constraint of pressure positivity into a set of equivalent linear constraints. It has been proven that the LDF OEDG method is PP if an “upwind” discrete Godunov–Powell source term is added solely to the evolution equations of the cell averages, under a condition achievable through a simple PP limiter. We have derived the PP CFL condition via a general convex decomposition of the cell averages and have also obtained the sharp PP CFL constraint through the optimal convex decomposition. The effectiveness of our method is validated through extensive one- and two-dimensional MHD tests. The numerical results demonstrate that the proposed method not only achieves the expected optimal convergence order but also exhibits robust performance in challenging MHD scenarios involving low density, low pressure, strong discontinuities, or low plasma-beta. Our future work includes extending the PP LDF OEDG schemes to numerical simulations of MHD on unstructured meshes.